Operation Summary for May 7, 2010 - May 10, 2010 |
The Integrated Luminosity for the period from 5/3/10 to 5/10/10 was 58.78 pb-1. Friday May 7 The day shift began with Operations monitoring the Recycler stash (184.7E10) and the Antiproton Source (Pbar) stack (6.9E10), with the Linac and operating normally, with the Main Injector (MI) waiting for beam, with the Tevatron (TeV) and Booster in access, with the NuMI, MiniBooNE, and Meson MTest experiments waiting for beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments in access. During the day shift, TeV personnel put the sector A0 abort kickers into conditioning mode. Operations established beam to Booster at 11:19 AM, with beam to Pbar and MiniBooNE by 11:35 AM, and with beam to NuMI by 12:02 PM. Water technicians reported finding and fixing two Meson magnets with LCW leaks. During the evening shift, TeV and D0 came out of access at 5:30 PM. Operations put the TeV through a dry squeeze at 7 PM. Operations conducted an orbit check and then put the TeV into HEP shot setup at 9 PM. Operations established store 7798 at 10:10 PM with an average initial luminosity (AIL) of 373.31E30. Booster personnel reported that they had swapped out a high voltage power supply for the notcher. Saturday May 8 The midnight shift began with Operations monitoring store 7798 (224.71E30) and the stash (86.8E10), with Pbar stacking (7.8E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, at 3:41 AM, Operations reported that the P1 transfer line corrector power supplies (VT703) output went too high and caused the line’s beam permit to trip off. Operators investigated and then loaded a backup file for the Pbar autotune program. A NuMI kicker (KPS6N) current went too low and caused the NuMI beam permit to trip off; it reset. Operations established beam to MTest at 4:30 AM. During the day shift, at 10:33 AM, Operations reported problems with the MI RF low and mid level front ends (FE). After consulting with experts, operators rebooted the FEs and MI resumed sending beam at 11:22 AM. Operations terminated store 7798 at 3:23 PM and put the TeV into shot setup at 3:34 PM. During the evening shift, Operations established store 7799 at 4:34 PM with an AIL of 400.6E30. Operations disabled the MTest beam at 10:01 PM. Sunday May 9 The midnight shift began with Operations monitoring store 7799 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations established beam to MTest at 4:22 AM. At 5:32 AM, Operations reported that the H- Source arc current was constantly alarming; operators adjusted the arc off timing. The Linac was working off the I- Source. During the day shift, Operations terminated store 7799 at 9 AM. Operators prepared D0 for access at 10:16 AM. Operations put the TeV into shot setup at 11:53 AM and established store 7802 at 12:58 PM with an AIL of 389.4E30. During the evening shift, at 5:17 PM, the TeV suffered a quench at sector D34. Operators and TeV personnel investigated, but could find no definite cause for the quench. Operations disabled the MTest beam at 6 PM. Operations reported at 7:17 PM that the Recycler missed an antiproton transfer due to a kicker (R:KPS2A) pre/fail to fire. Operations put the TeV through a dry squeeze at 8:47 PM. At 9:05 PM, Operations reported that Pbar lost an antiproton transfer when the stacktail didn’t turn off properly. Operators halted the transfers and resumed stacking. Operations put the TeV into shot setup at 10:09 PM and established store 7804 at 11:37 PM with an AIL of 332.33E30. Monday May 10 The midnight shift began with Operations monitoring store 7804 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations established beam to MTest at 4:35 AM. At 5:24 AM, Operations reported an A1 transfer line vacuum burst that caused a MI beam permit trip. An ion pump in TeV F sector had tripped off prior to the A1 line closure; everything reset. The Plans for Monday The plans for today are to stack, stash, and store. |
|
More Information Comments and Suggestions |
| last modified 05/10/2010 email Fermilab |
| Security, Privacy, Legal | |